OpenStudy (anonymous):

During hurricane season, Bob and Jen fill sand bags every weekend. On Saturday, they both worked together 4 hours and filled 32 sand bags. Sunday, Jen was feeling rather stiff and only worked 2 hours compared to Bob's 6 hours. They filled 36 sand bags on Sunday. How many sand bags can Jen fill in 6 hours?

OpenStudy (anonymous):

Let b=bob's rate of bags per hour; Let j=Jen's rate of bags per hour. So b+J=32/4 So bob and Jen can fill 8 bags in one hour when working together. Since bob works 6 hours and Jen works 2 hours, then bob's rate per hour times six hours and Jen's rate per hour times 2 hours equals 36 bags 6(b)+2(j)=36 bags Rearranging the first equation, we get b=8-j Substituting in to the second equation, we get 6(8-j)+2j=36 48-6j+2j=36 48-4j=36 48-36=4j 12=4j 3=j So 3 bags an hour is Jen's rate of filling bags of sand. Since the question asks how many bags Jen can fill in six hours, we will multiply six times 3 and get 6(3)=18 bags